# Revisiting Shooting Point Monte Carlo Methods for Transition Path Sampling

### v2

@@ II.1 Fixed- and Flexible-length Path Ensembles
Assuming two (meta)stable states A and B, we denote the ensemble of paths that connect these two states as $P_{\text{AB}}(X)$, where $X=\{x_{1},x_{2},...x_{L}\}$ is a discretized trajectory of length $L$ and each time slice or frame $x$ contains the coordinates (and possibly momenta) of all particles in the system.
The definition of $P_{\text{AB}}(X)$ depends on the desired path ensemble. For instance, a path distribution where all paths are of fixed length $L$ can be defined as
@@
Here, $H_{\text{AB}}(X)$ is unity if the path $X$ fulfills all requirements to be considered reactive and zero otherwise, as specified below.
Since pathways of different lengths have different measures, one must take special precautions to obtain a consistent definition of the transition path ensemble distribution. To do so, we consider, in fact, pathways that are of equal length $L_{\text{max}}$ and require that:
  1. 1.
    paths start in A, $x_{1}\in\text{A}$
  2. 2.
    paths reach B, i.e., there is an $i\leq L_{\text{max}}$ such that $x_{i}\in\text{B}$
  3. 3.
    all points of the path except $x_{1}$ are outside of A
If these conditions are met, $H_{\text{AB}}(X)$ is unity and it is zero otherwise. We denote the index of the first point in B as the path length $L(X)$. Note that this way of interpreting the distribution is only formal. In a practical path sampling scheme, one halts the integration of the equations of motion when a trajectory enters a stable state. Extending the trajectory (hypothetically) to full length $L_{\rm max}$ will then result in a valid pathway.
@@ II.2 Extended Space of Paths and Shooting Indices
In path sampling methods like aimless shooting and spring shooting, information on the previous shooting point is used to generate the new shooting point. To take this information into account, we denote with $k$ the index of a (configuration) point $x_{k}$ on the path $X$. Depending on the sampling scheme, this point can either be a shooting point directly or a point that is used to generate the next shooting point. We refer to $k$ as the shooting index in both cases since, even in the second case, the move is initiated from $k$. Common for both cases is that the index $k$, in some form or the other, evolves from trial to trial.
@@
that depends on the details of the generation algorithm. The new path and shooting point are then accepted with acceptance probability $p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}$. If $(X^{\text{n}},k^{\text{n}})$ is rejected, the old path and shooting point are counted again in the path or index ensembles and the corresponding averages.
To derive an acceptance probability for the Monte Carlo move, the joint probability density $P(X,k)$ must be specified. Since our goal is to sample the transition path ensemble $P_{\text{AB}}(X)$, we require that marginalizing $P(X,k)$ with respect to $k$ yields $P_{\text{AB}}(X)$,
$\sum_{k}P(X,k)=P_{\text{AB}}(X)\,.$ (6)
This condition is naturally fulfilled by defining $P(X,k)$ via the conditional probability $p(k|X)$, which has been described previously in the context of auxiliary variable methods Higdon (1998):
$P(X,k)=P_{\text{AB}}(X)p(k|X)\,.$ (7)
@@ II.3 Shooting Point Monte Carlo
The goal is now to construct a Markov chain Monte Carlo procedure for sampling the joint probability $P(X,k)=P_{\text{AB}}(X)p(k|X)$. In the subspace of the trajectories $X$ this procedure then samples the desired transition path ensemble $P_{\text{AB}}(X)$. Imposing detailed balance
@@
which can be satisfied with the Metropolis criterion. The generation probability can be further factorized as
@@
to separate the generation probabilities of the new path from the generation probabilities of the new shooting index. This factorization implies that first a new path is generated according to $p_{\text{gen}}(X^{\text{n}}|X^{\text{o}},k^{\text{o}})$ based on the old path and the old shooting point. Then, the new shooting point is generated according to $p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})$ based on the old shooting point as well as the old and new path. Another possibility, corresponding to a different factorization, is to first generate the new shooting point and then the new path or further factorize the shooting point generation process.
@@ II.4 Shooting Point Densities
$=\frac{1}{\langle L(X)\rangle_{X}}\int\text{d}X\,P_{\text{AB}}(X)\sum_{i=1}^{L(X)}\delta(x-x_{i})\,,$ (12)
where the delta function probes the occurrence of $x$ at position $x_{i}$ on a transition path $X$. In the above equation, the notation $\langle\cdots\rangle_{X}$ implies an average over the transition path ensemble $P_{\text{AB}}(X)$. Hence, $\langle L(X)\rangle_{X}$ is the average length of transition pathways. Note that shorter paths contribute fewer points to the distribution $\rho(x|\text{TP})$ defined in Eq. (II.4). For low-dimensional systems, this density can be computed numerically in a TPS simulation by histogramming points on the sampled transition pathways. In higher dimensions, it is more practical to consider the distribution $\rho(r|\text{TP})$ of a collective variable $r(x)$ for points on transition pathways. In either case, the distribution is a key indicator for the correct convergence of a path sampling scheme.
Figure 1: The standard shooting move is irreversible in the extended space framework. Starting from a path $X^{\text{o}}$ and previous shooting point $k^{\text{o}}$ (left), the shooting move is usually divided into two steps: first, a new shooting point $\hat{k}^{\text{o}}$ is picked and, second, the new path $X^{\text{n}}$ is generated from the selected point. On the new path the used shooting point has the index $k^{\text{n}}$. Although the reverse move (right) can recover the old path by selecting the same shooting index $k^{\text{n}}$ again, it is not possible to recover the old shooting index (indicated by the red arrow).
@@
For path sampling schemes that do not evolve the shooting point from trial to trial, each new shooting point can be generated independently by directly drawing it from $p(k|X)$, meaning that $p_{\text{sel}}(\hat{k}|X,k)=p(\hat{k}|X)$. In this case, the above expression reduces to
@@
To gain further insights into the relation of the attempted shooting point distributions in equations (II.4), (14) and  (15) to the extended space ensemble $P(X,k)$, we determine the distribution of shooting points, not attempted, but as obtained from $P(X,k)$ by integrating/summing over pathways $X$ and shooting indices $k$:
$\rho_{\text{ens}}(x)$ $=\int\text{d}X\,\sum_{k=1}^{L(X)}P(X,k)\delta(x-x_{k})$
@@
That means that $\rho_{\text{ens}}(x)$, the distribution of points after summing over $k$, is the same as the distribution of attempted shooting points one would obtain if shooting was performed by freshly drawing $k$ each trial from $p(k|X)$. Since $\rho_{\rm ens}(x)$ is independent of the specific shooting scheme yet depends on $p(k|X)$, we use it to assess the convergence of shooting schemes in the extended space $(X,k)$.
@@ II.5 One-way and Two-way TPS
We will next cast various TPS algorithms in the extended space formalism, starting with regular one-way and two-way shooting. In fixed- or flexible-length regular TPS, one draws the shooting index $k$ freshly at each round. As a result, there is no dependence of the picked shooting index on the previous shooting index $k^{\text{o}}$ and Markovianity is guaranteed without the extended space formalism. In fact, since regular TPS is carried out in trajectory space only, the acceptance rules derived for the extended space do not directly apply to this case. Vice versa, casting the algorithm unmodified to the extended space leads to irreversibility of the move (Fig. 1). However, we can construct an alternative TPS scheme that (a) is Markovian in the extended space, and (b) closely resembles the original two-way shooting algorithm.
We split the shooting move into three substeps as illustrated in Fig. 2. In the first part, the index shift, the shooting index is redrawn shifting the old shooting point $k^{\text{o}}$ to the point $\hat{k}^{\text{o}}$. In the second step, the shooting move itself is carried out from the point at $\hat{k}^{\text{o}}$. As the shooting move in most cases leads to a renumbering of points, we then define $\hat{k}^{\text{n}}$ as the index of the shooting point on the new path. As a last step, the shooting index is redrawn once again to ensure reversibility of the move, resulting in a new path and shooting index $(X^{\text{n}},k^{\text{n}})$.
Since we now work in the extended space, we need to specify the target shooting index probability $p(k|X)$ in order to obtain the full joint probability $P(X,k)$. To enforce a uniform shooting index distribution, we define $p(k|X)$ as:
@@
For both index shift moves (Fig. 2), we draw the shooting index from the uniform distribution to simplify the acceptance and keep the algorithm close to standard TPS, such that
@@
$p_{\text{sel}}(\hat{k}^{\text{n}}|X^{\text{n}},X^{\text{o}},\hat{k}^{\text{o}})=\frac{1}{L(X^{\text{n}})}\ \text{for}\ 1\leq k\leq L(X^{\text{n}})\,.$ (18)
Combining both shifts leads to a symmetric generation probability for the forward and the reverse move:
$p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})$ $=p_{\text{gen}}(k^{\text{o}}|X^{\text{o}},X^{\text{n}},k^{\text{n}})$
$=\frac{1}{L(X^{\text{n}})}\frac{1}{L(X^{\text{o}})}\,.$ (19)
After cancellation of the shooting index generation probabilities defined above, the acceptance probability based on Eqs. (II.3) and (17) is then given by
@@
In the analogous way, the shooting range algorithm Jung et al. (2017) can be formulated in the extended ensemble. In this case, shooting indices $k$ are only accepted in a range defined by a collective variable $r(x)$, corresponding to a shooting index distribution
$p(k|X)=\frac{h(r(x_{k}))}{\sum_{i}h(r(x_{i}))}\,.$ (21)
Here, $h(r)$ is unity if $a\leq r(x)\leq b$ with user defined bounds $a$ and $b$ and zero otherwise. The derived acceptance includes the above shooting index distribution and the condition must hold that all paths have at least one point within the shooting range Jung et al. (2017). Alternatively, a weight function $w$ can be defined for each $k$ to bias the shooting index selection
@@
Indeed, setting $w(r)=h(r)$ recovers eq (21). In this case, the acceptance criterion becomes
$p_{\text{acc}}\bigl{[}(X^{\text{o}},\hat{k}^{\text{o}})$ $\to(X^{\text{n}},\hat{k}^{\text{n}})\bigr{]}$
@@ II.6 Fixed-length Aimless Shooting
Aimless shooting with fixed path lengths is historically the basis for algorithms that evolve the shooting point from trial to trial. While there are different formulations of the algorithm in literature, we focus here on the two-point variant described in Ref. Peters (2010). Here, we start with an initial path $X^{\text{o}}$ of fixed path length $L$. The shooting index distribution is only nonzero at two points, $L/2-\Delta k$ and $L/2+\Delta k$, where $\Delta k$ is an adjustable parameter. This choice corresponds to the target shooting index distribution
@@
The two-point aimless shooting algorithm starts by selecting one of these two indices on the old path as the shooting index $\hat{k}^{\text{o}}$, which corresponds to the first index shift for regular TPS (Figure 2). Then, integration is performed in the backward direction by $L/2-\Delta k$ steps and in the forward direction by $L/2+\Delta k$ steps to obtain the new path $X^{\text{n}}$. This operation produces a new path on which the index of the common point, i.e., the index $\hat{k}^{\text{n}}$ of the point where the old and the new path intersect, can be either at $L/2-\Delta k$ or at $L/2+\Delta k$.
In principle, a second shift from $\hat{k}^{\text{n}}$ to $k^{\text{n}}$ would now be required to ensure that the move is reversible in the extended space. However, it is important to note that in fixed-length aimless shooting the shooting point does not evolve, in contrast to to its flexible-length counterpart. While there is a chance that the next shooting move is initiated from the same point as in the last trial, the shooting index is always drawn independently as $L/2-\Delta k$ or $L/2+\Delta k$ each trial, effectively erasing the memory of the previous shooting index. In other word, given a certain path $X^{\text{o}}$, determining the new shooting index is always done relative to the center $L/2$ of the path and does not require knowledge of the previous shooting index. As a result, the procedure is Markovian in path space and therefore does not require an extended space formalism with second shift. We introduce a second shift as for standard TPS, which allows us to apply the extended space formalism and infer the distribution of the shooting points from which a shooting is attempted via Eq. II.4:
@@
$p_{\text{acc}}\bigl{[}(X^{\text{o}},\hat{k}^{\text{o}})$ $\to(X^{\text{n}},\hat{k}^{\text{n}})\bigr{]}$

### v3

@@ II.1 Fixed- and Flexible-length Path Ensembles
Assuming two (meta)stable states A and B, we consider an ensemble of discretized trajectories $X=\{x_{1},x_{2},...x_{L}\}$ that connect A and B, where each time slice or frame $x$ contains the coordinates (and possibly momenta) of all particles in the system. We define a probability density in the ensemble of transition paths as $P_{\text{AB}}(X)$. For instance, a path distribution where all paths are of fixed length $L$ can be defined as
@@
Here, $H_{\text{AB}}(X)$ is unity if the path $X$ fulfills all requirements to be considered reactive and zero otherwise, as specified below, i.e. $x_{1}\in\mathrm{A},x_{L}\in\mathrm{B}$.
@@ II.2 Extended Space of Paths and Shooting Indices
In path sampling methods like aimless shooting and spring shooting, information on the previous shooting point is used to generate the new shooting point. To take this information into account, we denote with $k$ the index of a (configuration) point $x_{k}$ on the path $X$. Depending on the sampling scheme, this point can either be a shooting point directly or a point that is used to generate the next shooting point. The exact interpretation of $k$ will become clear later when we discuss several path sampling algorithms (see, for example, sections II.5, II.6, II.7 and II.8). We refer to $k$ as the shooting index in both cases since, even in the second case, the move is initiated from $k$. Common for both cases is that the index $k$, in some form or the other, evolves from trial to trial.
@@
that depends on the details of the generation algorithm. The new path and shooting index are then accepted with acceptance probability $p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}$. If $(X^{\text{n}},k^{\text{n}})$ is rejected, the old path and shooting index are counted again in the path or index ensembles and the corresponding averages.
To derive an acceptance probability for the Monte Carlo move, the joint probability density $P_{\mathrm{AB}}(X,k)$ must be specified. Since our goal is to sample the transition path ensemble $P_{\text{AB}}(X)$, we require that marginalizing $P_{\mathrm{AB}}(X,k)$ with respect to $k$ yields $P_{\text{AB}}(X)$,
$\sum_{k}P_{\mathrm{AB}}(X,k)=P_{\text{AB}}(X)\,.$ (6)
This condition is naturally fulfilled by defining $P_{\mathrm{AB}}(X,k)$ via the conditional probability $p(k|X)$, which has been described previously in the context of auxiliary variable methods Higdon (1998):
$P_{\mathrm{AB}}(X,k)=P_{\text{AB}}(X)p(k|X)\,.$ (7)
@@ II.3 Shooting Point Monte Carlo
The goal is now to construct a Markov chain Monte Carlo procedure for sampling the joint probability $P_{\mathrm{AB}}(X,k)=P_{\text{AB}}(X)p(k|X)$. In the subspace of the trajectories $X$ this procedure then samples the desired transition path ensemble $P_{\text{AB}}(X)$. Imposing detailed balance
@@
which can be satisfied with the Metropolis-Hastings criterion. The generation probability can be further factorized as
@@
to separate the generation probabilities of the new path from the generation probabilities of the new shooting index. This factorization implies that first a new path is generated according to $p_{\text{gen}}(X^{\text{n}}|X^{\text{o}},k^{\text{o}})$ based on the old path and the old shooting index. This probability includes all factors resulting from the propagation rules of the underlying dynamics Dellago et al. (2002); Bolhuis et al. (2002). Then, the new shooting index is generated according to $p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})$ based on the old shooting index as well as the old and new path. This factor describes the probability to select the new shooting index $k^{\text{n}}$. Another possibility, corresponding to a different factorization, is to first generate the new shooting index and then the new path or further factorize the shooting index generation process.
@@
Equation (II.3) can be satisfied by applying the Metropolis-Hastings criterion,
$p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}=$
$H_{\text{AB}}(X^{\text{n}})\min\biggl{[}1,\frac{p(k^{\text{n}}|X^{\text{n}})p_{\text{gen}}(k^{\text{o}}|X^{\text{o}},X^{\text{n}},k^{\text{n}})}{p(k^{\text{o}}|X^{\text{o}})p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})}\biggr{]}\,.$ (12)
@@ II.4 Shooting Point Densities
$=\frac{1}{\langle L(X)\rangle_{\mathrm{TP}}}\int\text{d}X\,P_{\text{AB}}(X)\sum_{i=1}^{L(X)}\delta(x-x_{i})\,,$ (13)
where the delta function probes the occurrence of $x$ at position $x_{i}$ on a transition path $X$. In the above equation, the notation $\langle\cdots\rangle_{\mathrm{TP}}$ implies an average over the transition path ensemble $P_{\text{AB}}(X)$. Hence, $\langle L(X)\rangle_{\mathrm{TP}}$ is the average length of transition pathways. Note that shorter paths contribute fewer points to the distribution $\rho(x|\text{TP})$ defined in Eq. (II.4). For low-dimensional systems, this density can be computed numerically in a TPS simulation by histogramming points on the sampled transition pathways. In higher dimensions, it is more practical to consider the distribution $\rho(r|\text{TP})$ of a collective variable $r(x)$ for points on transition pathways. In either case, the distribution is a key indicator for the correct convergence of a path sampling scheme.
Figure 1: The standard shooting move is irreversible in the extended space framework. Starting from a path $X^{\text{o}}$ and previous shooting index $k^{\text{o}}$ (left), the shooting move is usually divided into two steps: first, a new index $\hat{k}^{\text{o}}$ is picked and, second, the new path $X^{\text{n}}$ is generated from the selected point. On the new path the used shooting point has the index $k^{\text{n}}$. Although the reverse move (right) can recover the old path by selecting the same shooting index $k^{\text{n}}$ again, it is not possible to recover the old shooting index (indicated by the red arrow).
@@
For path sampling schemes that do not evolve the shooting index from trial to trial, each new shooting index can be generated independently by directly drawing it from $p(k|X)$, implying that $p_{\text{sel}}(\hat{k}|X,k)=p(\hat{k}|X)$. In this case, the above expression reduces to
×$\times\sum_{\hat{k}=1}^{L(X)}p(\hat{k}|X)\delta(x-x_{\hat{k}})\biggr{]}\,,$ (15)
where the sum over $k$ cancels and, renaming $\hat{k}$ as $k$, we arrive at:
@@
We can then relate the attempted shooting point distributions in Eqs. (II.4), (16) and  (17) to the extended space ensemble $P_{\mathrm{AB}}(X,k)$. For that, we determine the distribution of shooting points, not attempted, but as obtained from $P_{\mathrm{AB}}(X,k)$ by integrating/summing over pathways $X$ and shooting indices $k$:
$\rho_{\text{ens}}(x)$ $=\int\text{d}X\,\sum_{k=1}^{L(X)}P_{\mathrm{AB}}(X,k)\delta(x-x_{k})$
@@
This means that $\rho_{\text{ens}}(x)$, the distribution of points after summing over $k$, is the same as the distribution of attempted shooting points (Eq. (16)) one would obtain if shooting was performed by freshly drawing $k$ each trial from $p(k|X)$. Since $\rho_{\rm ens}(x)$ is independent of the specific shooting scheme yet depends on $p(k|X)$, we use it to assess the convergence of shooting schemes in the extended space $(X,k)$ as shown in Figs. 5 and 6.
@@ II.5 One-way and Two-way TPS
We will next cast various TPS algorithms in the extended space formalism, starting with regular one-way and two-way shooting. In fixed- or flexible-length regular TPS, one draws the shooting index $k$ freshly at each round. As a result, there is no dependence of the picked shooting index on the previous shooting index $k^{\text{o}}$. In fact, since regular TPS is carried out in trajectory space only, the acceptance rules derived for the extended space do not directly apply to this case. Vice versa, casting the algorithm without modification to the extended space leads to irreversibility of the move (Fig. 1). However, we can construct an alternative TPS scheme that is reversible in the extended space, and closely resembles the original two-way shooting algorithm.
We split the shooting move into three substeps as illustrated in Fig. 2. In the first part, the index shift, the shooting index is redrawn shifting the old shooting index $k^{\text{o}}$ to the point $\hat{k}^{\text{o}}$. In the second step, the shooting move itself is carried out from the point at $\hat{k}^{\text{o}}$ leading to the new path $X^{\text{n}}$. As the shooting move in most cases leads to a renumbering of points, we then define $\hat{k}^{\text{n}}$ as the index of the shooting point on the new path. As a last step, the shooting index is redrawn once again to ensure reversibility of the move. The whole shooting move results in a new path and shooting index $(X^{\text{n}},k^{\text{n}})$.
Since we now work in the extended space, we need to specify the target shooting index probability $p(k|X)$ in order to obtain the full joint probability $P_{\mathrm{AB}}(X,k)$. To enforce a uniform shooting index distribution, we define $p(k|X)$ as:
@@
For both index shift moves (steps (2) and (4) in Fig. 2), we draw the shooting index from the uniform distribution to simplify the acceptance and keep the algorithm close to standard TPS, such that
@@
$p_{\text{sel}}(k^{\text{n}}|X^{\text{n}})=\frac{1}{L(X^{\text{n}})}\ \text{for}\ 1\leq k\leq L(X^{\text{n}})\,.$ (20)
Combining both selection probabilities leads to a symmetric index generation probability for the forward and the reverse move, $p_{\text{gen}}^{\mathrm{fw}}=p_{\text{gen}}^{\mathrm{rv}}=p_{\text{sel}}(\hat{k}^{\text{o}}|X^{\text{o}})p_{\text{sel}}(k^{\text{n}}|X^{\text{n}})$. Therefore the acceptance probability based on Eqs. (II.3) and (19) is then given by
@@
In the analogous way, the shooting range algorithm Jung et al. (2017) can be formulated in the extended ensemble. In this case, shooting indices $k$ are accepted according to a weighting function $w(r(x_{k}))$ along a collective variable $r(x)$, corresponding to a shooting index distribution:
@@
In this case, the acceptance criterion becomes
@@
Setting $w(r)$ to unity if $a\leq r(x)\leq b$ with user defined bounds $a$ and $b$ and zero otherwise results in the original shooting range algorithm Jung et al. (2017).
@@ II.6 Fixed-length Aimless Shooting
Aimless shooting with fixed path lengths is historically the basis for algorithms that evolve the shooting point from trial to trial. While there are different formulations of the algorithm in literature, we focus here on the two-point variant described in Ref. Peters (2010). Here, we start with an initial path $X^{\text{o}}$ of fixed path length $L$. We assume an even path length in the following section, for odd path lengths integration steps need to be adjusted accordingly. The shooting index distribution is only nonzero at two points, $L/2-\Delta k$ and $L/2+\Delta k$, where $\Delta k$ is an adjustable parameter. This choice corresponds to the target shooting index distribution
@@
The two-point aimless shooting algorithm starts by selecting one of these two indices on the old path as the shooting index $\hat{k}^{\text{o}}$, which corresponds to the first index shift for regular TPS (Figure 2). For a positive shift, integration is performed in the backward direction by $L/2+\Delta k-1$ steps and in the forward direction by $L/2-\Delta k$ steps to obtain the new path $X^{\text{n}}$ (vice versa for a backward shift). This operation produces a new path on which the index of the common point, i.e., the index $\hat{k}^{\text{n}}$ of the point where the old and the new path intersect, can be either at $L/2-\Delta k$ or at $L/2+\Delta k$.
In principle, a second shift from $\hat{k}^{\text{n}}$ to $k^{\text{n}}$ would now be required to ensure that the move is reversible in the extended space. However, it is important to note that in fixed-length aimless shooting the shooting index does not evolve, in contrast to to its flexible-length counterpart. While there is a chance that the next shooting move is initiated from the same point as in the last trial, the shooting index is always drawn independently as $L/2-\Delta k$ or $L/2+\Delta k$ each trial, effectively erasing the memory of the previous shooting index. In other word, given a certain path $X^{\text{o}}$, determining the new shooting index is always done relative to the center $L/2$ of the path and does not require knowledge of the previous shooting index. As a result, the procedure is Markovian in path space and therefore does not require an extended space formalism with second shift. We introduce a second shift as for standard TPS, which allows us to apply the extended space formalism and infer the distribution of the shooting points from which a shooting is attempted via Eq. II.4: